40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 174 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  8405   Wed Apr 3 18:22:00 2013 JenneUpdateElectronicsPOP110 re-implemented

I have re-implemented POP110.  The cable coming from the AS110 diode is disconnected, labeled, and sitting in the cable tray next to the LSC rack. 

Now the POP diode path is:

 

Thorlabs 10CF ----many meters of heliax cable-----> Bias Tee ------> RF amplifier ------> Splitter ------> Bandpass 21.7MHz --------> POP22 demod board

                                                                                                                   |                                                                                    |

                                                                                                                   |                                                                                    |

                                                                                                                   V                                                                                  V

                                                                                                            POP DC                                                                        High pass 100MHz

                                                                                                                                                                                                        |

                                                                                                                                                                                                        |

                                                                                                                                                                                                       V

                                                                                                                                                                                              Lowpass 150MHz

                                                                                                                                                                                                       |

                                                                                                                                                                                                       |

                                                                                                                                                                                                      V

                                                                                                                                                                                        POP110 demod board

  8404   Wed Apr 3 17:40:18 2013 JamieConfigurationElectronicsputting together a 110 MHz LSC demod board

I started to look into putting together a 110 MHz demod board to be used as POP110 (see #8399).

We have five spare old-skool EuroCard demod boards (LIGO-D990511).  From what I gather (see #4538, #4708) there are two modifications we do to these boards to make them ready for prime time:

  • appropriate LP filter at PD RF input (U5 -> MC SCLF-*)
  • swap out T1 transformer network with a commercial phase shifting power splitter (MC PQW/PSCQ)

#4538 also describes some other modifications but I'm not sure if those were actually implemented or not:

  • removal of the attenuator/DC block/ERA-5 amp sections at the I/Q outputs
  • swap ERA-5 amp with "Cougar"(?) amp at LO input.

What we'll need for a 110 demod:

I'll scrounge or order.

  8403   Wed Apr 3 16:03:59 2013 ChloeUpdate QPD Voltage Regulators

The voltage regulator on the QPD breadboard seems to be having problems... yesterday Eric helped me debug my circuit and discovered that the +12V regulator was overheating, so we replaced it. Today, I found that the -12V regulator was also doing the same thing, so I replaced it. However, it's still overheating. We checked all of the setup for the power regulators yesterday, so I'm not sure what's wrong.

I've also noticed that not all the connections on the breadboard that I've been using seem to work - I may search for a new breadboard in this case. Need to check I'm not doing something stupid with that.

  8402   Wed Apr 3 15:00:24 2013 JamieSummaryElectronicsSorensen supplies in LSC rack (1Y2)

I investigated the situation of the two Sorensen supplies in the LSC rack (1Y2).  They are there solely to supply power to the LSC LO RF distribution box.  One is +18 V and the other is +28 V.  All we need to do is make a new longer cable with the appropriate plug on one end (see below), long enough to go from the bottom of the 1Y3 rack to the top of 1Y2, and we could move them over quickly.  Some sort of non-standard circular socket connector is used on the distribution box:

20130403_141842.jpg

It could probably use thicker conduction wire as well.

If someone else makes the cable I'll move everything over.

  8401   Wed Apr 3 14:46:17 2013 GabrieleSummaryLSCError signal simulation in PRMI

Here is a summary of a simulation of the error signal behavior in the PRMI configuration. The main parameters are:
L_PRC = 6.7538 m
Schnup = 0.0342 m
fmod1 = 11.065399e6 Hz
fmod2 = 5 * fmod1

prmi_prclsweep_pop22.pngprmi_prclsweep_pop110.png

These two plots shows the response of the POP22 and POP110 signals (in almost arbitrary units) to a PRCL sweep around the resonance. The splitting of the 55 sideband peaks is well visible in the second plot. It is due to the fact that the 55MHz sidebands are not perfectly matched to the PRC length

prmi_michsweep_pop22.pngprmi_michsweep_pop110.png

The same thing when sweeping MICH. The peaks are wider and it is not possible to see the splitting.

prmi_prclsweep_errsig_not_tunedphase.pngprmi_michsweep_errsig_not_tunedphase.png

These are the error signals (REFL11_I/Q and REFL_55_I/Q) as a function of the PRCL (left) and MICH (right) sweep. Here the demodulation phases are not properly tuned. This is just to show that when the phase is wrong, you can get multiple zero crossings (in this case only in the Q signals, but in general also in I) close to resonance.

prmi_prclsweep_errsig_tunedphase.pngprmi_michsweep_errsig_tunedphase.png

If the phases are tuned in order to maximize the slope of the I signals with respect to PRCL, one gets these "optimized phase" responses. It is that the phase does not correspond to the one that makes the PRCL peak to peak signal small in Q. The Q signals are indeed flat around resonance for a PRCL motion, but they deviate quite a lot from zero when moving more far from resonance. Moreover, both the REFL_55 error signals (I for PRCL and Q for MICH) are crossing again zero at two additional positions, but those are quite far from the resonance point.

prmi_prclsweep_triggering.pngprmi_michsweep_triggering.png

These plots just show the PRCL and MICH error signals together with the POP22 and POP110 signals, to give an idea of the level of triggering that might be needed to be inside the linear range. It seems that if we trigger on POP22 when using the REFL55 signal we loose a bit of linear range, but not that much.

prmi_prclsweep_linearized_signals.pngprmi_michsweep_linearized_signals.png

If you reached this point it means you're really interested in this topic, or maybe you have nothing better to do... However, this plot shows the effect of linearization of the error signal, obtained dividing them by the proper POP22/110 signal. The linear range is increased, but unfortunately for the 55 signals, the additional zero crossing I was mentioning before creates two sharp features. Those are however quite outside the triggering region, so they should not be harmful.

  8400   Wed Apr 3 14:45:34 2013 JamieUpdateComputersupdated EPICS database (channels selected for saving)

Quote:

I modified /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini to include the C1:LSC-DegreeOfFreedom_TRIG_MON channels.  These are the same channel that cause the LSC screen trigger indicators to light up. 

I vaguely followed Koji's directions in elog 5991, although I didn't add new grecords, since these channels are already included in the .db file as a result of EpicsOut blocks in the simulink model.  So really, I only did Step 2.  I still need to restart the framebuilder, but locking (attempt at locking) is happening.

The idea here is that we should be able to search through this channel, and when we get a trigger, we can go back and plot useful signals (PDs, error signals, cotrol signals,....), and try to figure out why we're losing lock. 

Rana tells me that this is similar to an old LockAcq script that would run DTT and get data.

EDIT:  I restarted the daqd on the fb, and I now see the channel in dataviewer, but I can only get live data, no past data, even though it says that it is (16,float).  Here's what Dataviewer is telling me:

Connecting to NDS Server fb (TCP port 8088)
Connecting.... done
read(); errno=0
LONG: DataRead = -1
No data found

read(); errno=9
read(); errno=9
T0=13-03-29-08-59-43; Length=432010 (s)
No data output.
 

I seem to be able to retrieve these channels ok from the past:

controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ tconvert 1049050000
Apr 03 2013 18:46:24 UTC
controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ ./general/getdata -s 1049050000 -d 10 --noplot C1:LSC-PRCL_TRIG_MON
Connecting to server fb:8088 ...
nds_logging_init: Entrynds_logging_init: Exit
fetching... 1049050000.0
Hit any key to exit: 
controls@pianosa:/opt/rtcds/caltech/c1/scripts 0$ 

Maybe DTT just needed to be reloaded/restarted?

  8399   Wed Apr 3 03:18:23 2013 JenneUpdateLockingPRMI locked with REFL55 I&Q for more than 1 minute

[Rana, Gabriele, Jenne]

We have now locked the PRMI using REFL55 I&Q for more than one minute!!!!!

PRMI_sb_80sec_2Apr2013.png

This isn't really the most useful plot as is, but it was created using:

/opt/rtcds/caltech/c1/scripts/general/getdata C1:LSC-POP22_I_ERR_DQ C1:LSC-REFL55_I_ERR_DQ C1:LSC-REFL55_Q_ERR_DQ C1:LSC-MICH_IN1_DQ C1:LSC-MICH_OUT_DQ C1:LSC-PRCL_IN1_DQ C1:LSC-PRCL_OUT_DQ -d 80 -s 1049013520 -c

This is just one of several long lock stretches.  If I can get the TRIG_MON channels to be saved, we can automatically (versus my by-hand search) find lock stretches and make this kind of plot.  Although we want them saved in some raw format so we can zoom in on selected axes, I think.  This might require some python-fu from Jamie, or learning of python-fu for Jenne.

The secret sauce:

* The big key was changing REFL55's phase.  It was -4 when we looked at the I&Q signals, and minimized the PRCL information in the Q-phase.  We were able to get short lock stretches with this.  During these stretches, Rana changed the REFL55 phase until the lock sounded (audibly) quieter.  The final phase we settled on was +26. As we changed the phase, the lock stretches got longer and longer. 

* We also tweaked up the POP22 phase.  It was close from our previous efforts of looking at non-locked time series, but we perfected it by minimizing the signal in the Q-phase during lock stretches.  We also found that it drifted (according to this method) by ~5 degrees over ~half an hour (I don't remember the exact time between our phase tunings).

* POP22's low pass filters (both options, ELP10 and ELP50) must be OFF for any lock to be acquired.  Turning on either filter prevents locking.

* Normalization helped a lot.  Without normalization we weren't really able to catch any locks, certainly not of any significant length. (0.004, using POP22I, for both MICH and PRCL). 

* Parameters:

       ** Normalization:  use POP22I for both MICH and PRCL, value = 0.004

       ** Input matrix:  MICH with REFL55Q, value = 0.01; PRCL with REFL55I, value = 0.01  (we used the small number in the matrix so our servo gains weren't too tiny).

       ** POP22 lowpass filters OFF

       ** Analog whitening OFF for REFL55, POP22.

       ** Analog gain for REFL55 I&Q = 27 dB

       ** Analog gain for POP22 I&Q = 15 dB

       ** Output matrix:  MICH with -1 to ITMX, +1 to ITMY.  PRCL with +1 to PRM.

       ** Servo gains:  PRCL = 0.75;  MICH anywhere between -3 and -20.  Best in the -8 to -15 range.

       ** Vio2 filters in ITMX, ITMY, PRM (all actuated-on mirrors) were OFF. (Still need to lower the Q on these so they don't ring).

      ** PRCL and MICH triggering on POP22I.  The trigger-off was always 20, but the trigger-on changed throughout the night from ~170 to ~50.  I think 130 was a trigger value for at least some of the long-time locks.

      ** Low frequency seismic was small (i.e. no anomalous 0.1 Hz - 1 Hz noise) during successful lock times.  (Not to say it must be low, but it was low when we were able to lock for long stretches).

Things we had looked at and thought about throughout the evening:

* Oplev calibration.  See elog 8391 and 8393.  Optimized BS and PRM to reduce yaw angular motion.

* Actuators all functioning as expected.  We checked transfer functions of MICH_OUT/MICH_IN1 for locking with different optics, to ensure that at high frequency the response was 1/f^2.  Also, we locked MICH with (a) both ITMs, (b) BS, (c) ITMX and (d) ITMY.  We locked the PR-ITMY half-cav with (a) PRM and (b) ITMY.  We locked the PR-ITMX half-cav with (a) PRM and (b) ITMX.  Thus, we conclude that all of the PRMI-related optics are functioning as expected.

* Realigned REFL55 beam onto PD.  It was clipping a bit, so the DC power wasn't steady (when ITMs were misaligned, PRM aligned).  After alignment, the DC power as seen on a 'scope was much smoother.

* Turning off the limiters for the MICH and PRCL control signals allowed us to hear a high-pitched whine.  From looking at the time series, it's predominantly in MICH_OUT.  Rana speculates that perhaps the normalization is causing the UGF to wander temporarily to an unstable place.  For a time there was a high-Q peak between 500 and 600Hz, but reducing the gain (of MICH?) eliminated that.  Then we heard several times, irrespective of gain setting, the ~400Hz broad peak (I say broad because I was able to see it on DTT looking at the error and control signals, and it spanned +/-100Hz).

Things to investigate:

* Is there a good reason that we should switch to triggering on POP110, rather than the current POP22?  From Gabriele, Jamie and my Finesee/Mist modelling last week, without the arms, the 11MHz and 55MHz resonate at different PRC lengths.  If this difference is very small, then we are fine, but if the difference is large, it could be causing trouble - we're trying to catch the lock at the linear part of the 55MHz signal, but if that does not coincide with the linear part of the 11MHz signal, we're doing the wrong thing.  

* For the POP normalization, should we be using the amplitude or the power ( POP22 or sqrt(POP22) )?  Why?  Look at this with a modelling sweep and/or analytically.

* Look at different noise sources, potentially sensing noise, coil actuator noise,.....  We should check these out, and make sure we're not limited by anything obvious. 

Other To-Do:

* Make a "restore" medm screen, rather than restore script.  IFO Configure restore script can pull in values from the screen (EPICS values).  One screen per configuration. 

* Get TRIG_MON signals saved, write script to search for triggered lock times (between given gps times), then plot interesting signals for just before lock, during lock, and until just after a lockloss.

  8398   Wed Apr 3 01:32:04 2013 JenneUpdateComputersupdated EPICS database (channels selected for saving)

I modified /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini to include the C1:LSC-DegreeOfFreedom_TRIG_MON channels.  These are the same channel that cause the LSC screen trigger indicators to light up. 

I vaguely followed Koji's directions in elog 5991, although I didn't add new grecords, since these channels are already included in the .db file as a result of EpicsOut blocks in the simulink model.  So really, I only did Step 2.  I still need to restart the framebuilder, but locking (attempt at locking) is happening.

The idea here is that we should be able to search through this channel, and when we get a trigger, we can go back and plot useful signals (PDs, error signals, cotrol signals,....), and try to figure out why we're losing lock. 

Rana tells me that this is similar to an old LockAcq script that would run DTT and get data.

EDIT:  I restarted the daqd on the fb, and I now see the channel in dataviewer, but I can only get live data, no past data, even though it says that it is (16,float).  Here's what Dataviewer is telling me:

Connecting to NDS Server fb (TCP port 8088)
Connecting.... done
read(); errno=0
LONG: DataRead = -1
No data found

read(); errno=9
read(); errno=9
T0=13-03-29-08-59-43; Length=432010 (s)
No data output.

 

  8397   Tue Apr 2 23:14:02 2013 ranaUpdateLSC22/110MHz path for POP

Quote:

We could not find a power supply slot for the amplifiers on the LSC rack. We had to put a temporary power supply in contradiction to our 'no temporary power supply' policy.

 After 1 month, its hard to imagine that this could not have been fixed by putting in a proper fuse and fuse block. I will remove this tomorrow if I still find it this way in the bottom of the rack.

There are also 2 Sorensen switching supplies in the bottom of the LSC rack (with all of our sensitive demod boards). These should also be moved over to the old 'digital' LSC rack tomorrow for the post meeting lab cleanup.

Use fuse blocks with fuses with appropriate ampacity.

  8396   Tue Apr 2 22:39:17 2013 AnnalisaUpdateAuxiliary lockingBeat note between "Alberto" NPRO laser and PSL laser

 

 The beat note between the PSL laser and the "Alberto" NPRO laser has been measured. In particular, for each PSL temperature, more than one Aux laser frequency has been found.

The second of the three curves seems to be more stable than the other two, even if a "step" trend can be found in all of them (maybe due to the frequency change of the NPRO laser as a function of the crystal temperature continuous tuning, as mentioned in the previous elog). This is the reason why the points are not perfectly aligned, and the errors on the fit parameters are so big.

 

 

  8395   Tue Apr 2 21:11:42 2013 RanaUpdateoptical tablesOptical Table Toolboxes Update

Quote:

A heavy duty plastic box is the likeliest candidate for the optical table toolbox. It measures 5 9/16 in. x 11 5/8 in. x 4 5/8 in. and fits all the tools comfortably. ( http://www.mcmaster.com/#plastic-bin-boxes/=m4yh4m  ,  under Heavy Duty Plastic Bin Boxes)

The list of tools has been updated to include a pen and a wire cutter as well as everything previously stated.

In addition, Steve has recommended that boxes should be secured to the walls or surfaces near the optical tables as opposed to the optical tables themselves, as to keep the tables from wobbling when tools are being exchanged.

A diagram of tentative box placements will go out soon.

 No, the small boxes must be attached to the optical tables. They won't be heavy enough to change the table tilt.

Also, all tools must be color coded according to the optical table using the 3M Vinyl table color code:

http://www.3m.com/product/images/Vinyl-Electrical-Color-Tape-300.jpg

  8394   Tue Apr 2 20:52:35 2013 ranaUpdateComputersiMac bashed

 I changed the default shell on our control room iMac to bash. Since we're really, really using bash as the shell for LIGO, we might as well get used to it. As we do this for the workstations, some things will fail, but we can adopt Jamie's private .bashrc to get started and then fix it up later.

  8393   Tue Apr 2 18:19:30 2013 JenneUpdateSUSBS, PRM oplev servos improved

 

[Gabriele, Jenne]

We have implemented 4Hz resonant gains for both PRM and BS yaw.  The filter was already in place for PRM Yaw, so we just turned it on, but we also copied the filter over to BS Yaw.  We also changed the 3.3Hz res gain and the ELP for the PRM servo to match the BS servo, since after implementing the 4Hz gain, PRM was still much noisier than BS.  Now the 2 servos match, and PRM is a little quieter.  We hope that tonight's locking might be a little more stable after this work.

PRM_servo_matches_BS.pdf

prm_bs_optical_levers_comparison.pdf

  8392   Tue Apr 2 17:23:23 2013 SteveUpdate40m Upgradingnew ETMY optical table & enclosure in place

 Enclosure is at the east end. It has it's bottom o-ring in place. It will be ready for optics tomorrow around 5pm

I have to shim out the enclosure, finish leveling the table and cut surgical tubing O-ring for the top.

 

  8391   Tue Apr 2 16:55:34 2013 JenneUpdateSUSoplev calibration online for ITMs, BS, PRM

[Jenne, Gabriele]

We have put in a new EPICS input into the SUS library part, just before the OL_PIT and OL_YAW filter banks, so that the IN1 point is calibrated to microradians.  I recompiled all SUS-related models.  The OPTLEV_SERVO screen has been changed, so that you can see the calibration, and enter a value.  The gains have been reduced by a factor reciprocal to the calibration, so the loop gain is the same.

ETMs, SRM and MCs all have "calibration" numbers of 1, so the numbers aren't really calibrated, they're just the same as always.

It looks like the PRM and the BS are moving significantly (factor of ~30) more than the ITMs at a few Hz!  (Y-axis of plot is urad/rtHz)

comparison_opl.pdf

EDIT JCD:  We need to fix up the MEDM QPD indicators, and the OpLev red lights on the watchdog screen, so they match the new numbers.  Also, Rana turned on the output limiters to 2000 for all oplev servos.

  8390   Tue Apr 2 16:13:10 2013 ranaUpdateSUSoplev calibration for ITMX, BS

  Very good - now you need to just put the cal factor into the filter banks so that the PERROR and YERROR signals are in microradians all the time.

 

EDIT JCD:  In progress.

  8389   Tue Apr 2 15:58:40 2013 JenneUpdateSUSoplev calibration for ITMX, BS

[Jenne, Gabriele]

Optical lever calibrations:

ITMX pit calibration = -9.07 cts/mrad

ITMX yaw calibration = -12.33 cts/mrad

ITMX plot:opl_itmx.png

BS pit calibration = -22.86 cts/mrad

BS yaw calibration = -24.14 cts/mrad

BS plot:opl_bs.png

 

Method:  Similar to Manasa and Yuta's method last month.  We mounted each oplev QPD on a micrometer translation stage, centered the beam using the steering mirror, then used tdsavg to get 10 second averages of the _INMON channel for various settings of the micrometer stage.  For BS, we had to take out the PRM oplev to make room for the translation stage.  All QPDs were remounted in their original positions, within less than 1mm.  Measured the out-of-vac distances with the laser disto-meter, and the invac distances from the optic to the window from the CAD drawing.

Copying from other elog entries,

elog 8232:

We calibrated oplev for ITMY. Calibration factor for C1:SUS-ITMY_OL(PIT|YAW)_IN1 are;
  OLPIT: 6.29 +/- 0.11 counts/mrad
 
OLYAW: 5.74 +/- 0.09 counts/mrad

 

elog 8221:

We calibrated oplev for PRM. Calibration factor for C1:SUS-PRM_OL(PIT|YAW)_IN1 are;
  OLPIT: 15.6 +/- 0.3 counts/mrad
  OLYAW: 17.8 +/- 0.3 counts/mrad

  8388   Tue Apr 2 11:08:16 2013 KojiUpdateLSCPRMI lock

> The two REFL55 signals

Wow! It's a good news.
I think this is our first ever lock of PRMI with the REFL I/Q signals.

We kept having difficulty to obtain MICH from the REFL beam.

Next time could you make calibration of REFL55 MICH and AS55 MICH and compare the ratio with any simulation?

  8387   Tue Apr 2 10:22:37 2013 Rana, Gabriele, JenneUpdateLSCPRMI lock

We locked the PRMI, this time really on the sidebands, using the two REFL55 signals.

Here are the parameters: triggering on POP22_I in at 140, out at 20. No normalization. MICH gain -0.15, PRCL gain 0.1

It seems that the lock is not very stable. It seems likely to come from some large angular motion of one of the mirrors. We'll need to calibrate the optical lever signals to understand which one is moving too much.

 

  8386   Mon Apr 1 23:22:17 2013 AnnalisaUpdateAuxiliary lockingBeat note between "Alberto" NPRO laser and PSL laser

I measured the beat note between the "Alberto" NPRO laser and the PSL varying the PSL temperature and find the matching NPRO temperature that gave the beat.

I first switched off the FSS loop for the PSL, then I varied its temperature and switched on the loop back.

PSL temperature has been varied starting from 31.88 °C (its starting temperature) down to 23.88 by 1°C step, and then from 31.88 °C up to 36.92 °C, always with a 1°C step.

For each PSL temperature, the NPRO temperature was varied as well, in way to find the temperature to have a beat note between the two.

The trend of the NPRO laser temperature reminds the frequency change of the laser as a function of the crystal temperature continuous tuning.

I made measurements only for the first temperature of the NPRO laser which gave me the beat note. Tomorrow I'm going to find the beat note also for higher frequencies of the NPRO laser.

 

  8385   Mon Apr 1 17:59:04 2013 JenneUpdateLockingPRMI videos

[Gabriele's work, I'm just spectating]

Annalisa is working on finding the PSL/AUX laser beatnote, so the PSL temp is changing, but Gabriele is still able to lock.  Here are some videos:

 

 

  8384   Mon Apr 1 16:35:42 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

Quote:

A heavy duty plastic box is the likeliest candidate for the optical table toolbox. It measures 5 9/16 in. x 11 5/8 in. x 4 5/8 in. and fits all the tools comfortably. ( http://www.mcmaster.com/#plastic-bin-boxes/=m4yh4m  ,  under Heavy Duty Plastic Bin Boxes)

The list of tools has been updated to include a pen and a wire cutter as well as everything previously stated.

In addition, Steve has recommended that boxes should be secured to the walls or surfaces near the optical tables as opposed to the optical tables themselves, as to keep the tables from wobbling when tools are being exchanged.

A diagram of tentative box placements will go out soon.

 I also took every allen key I can find so they can be sorted. They will be back in the appropriate drawer locations soon.

  8383   Mon Apr 1 16:24:09 2013 JamieFrogsLSCPD whitening switching fixed (loose connection at break-out box)

Quote:

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

This turned out to just be a loose connection of the ribbon cable from Contec board in the LSC IO chassis at the BIO break-out box.  The DSUB connector at the break-out box was not strain relieved!  I reseated the connector and strain relieved it and now everything is switching fine.

20130401_161917.jpg

20130401_161858.jpg

I wonder if we'll ever learn to strain relieve...

  8382   Mon Apr 1 16:16:16 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

A heavy duty plastic box is the likeliest candidate for the optical table toolbox. It measures 5 9/16 in. x 11 5/8 in. x 4 5/8 in. and fits all the tools comfortably. ( http://www.mcmaster.com/#plastic-bin-boxes/=m4yh4m  ,  under Heavy Duty Plastic Bin Boxes)

The list of tools has been updated to include a pen and a wire cutter as well as everything previously stated.

In addition, Steve has recommended that boxes should be secured to the walls or surfaces near the optical tables as opposed to the optical tables themselves, as to keep the tables from wobbling when tools are being exchanged.

A diagram of tentative box placements will go out soon.

  8381   Mon Apr 1 16:13:50 2013 ChloeUpdate QPD

Because we would like to get started on testing mount vibrations as soon as possible, I've been trying to get one of the other QPDs we found to work with the summing/subtracting circuit on a breadboard. I've been using a power supply that I think Jamie built 15 years ago... which seems to be broken as of today, since I no longer read any signal from it with an oscilloscope.

I tried using a different power supply, but I still can't read any change in signal with the QPD for any of the quadrants when using a laser pointer to shine light on it. I'll be working with Eric on this later this week. In the meantime, I'll try and come up with a shopping list for the nicer QPD circuit that'll be a longer term side project.

  8380   Mon Apr 1 09:25:35 2013 JenneMetaphysicsLSCLock of PRMI on sidebands

[Gabriele, Jenne]

I put a notch in FM10 for both MICH and PRCL at 628Hz, to try to prevent us from exciting the mode that Gabriele saw on Friday.  Since those filter banks were all full, I have removed an ELP50 (ellip("LowPass",4,1,40,50)).  I write it down here, so we can put it back if so desired.

  8379   Mon Apr 1 09:05:09 2013 Jenne, GabrieleConfigurationLSCPOP22 configuration

On Friday we modified the POP22 set up: now the PD output goes to a bias tee. The DC output goes to the ADC board, while the RF output goes to an amplifier (Mini-circuits ZFL-1000LN+), to a band pass filter at 21.4 MHz and then to the ADC

  8378   Sun Mar 31 17:26:32 2013 ranaUpdateLSCLock of PRMI on sidebands

 

 Our first move has to be fixing the whitening switching for REFL55. That's the configuration we need to start and then move onto REFL165 to get to FPPRMI.

  8377   Fri Mar 29 19:58:24 2013 Gabriele, JenneFrogsLSCAnalog whitening filter of REFL55 not switching

Quote:

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

The digital one is actually switching. We decided to keep the digital de-whitening on to compensate for the analog one. Otherwise we get a very bad shape of the PDH signal. Sorry Rana...

 I forgot to say that the analog gain of the REFL55 channels has been reduced to 9db

  8376   Fri Mar 29 19:56:02 2013 GabrieleMetaphysicsLSCLock of PRMI on sidebands

I finally managed to get long stretches of PRMI lock, up to many minutes. The lock is not yest very stable, it seems to me that we are limited by some yaw oscillation that I could not trace down. The oscillation is very well visible on POP.

Presently, PRCL is controlled with REFL55_I, while MICH is controlled with AS55_Q. This configuration is maybe not optimal from the point of view of phase noise couplings, but at least it works quite well. I believe that the limit on the length of locks is given by the angular oscillation. I attach to this entry few plots showing some of the lock stretches. The alignment is not optimal, as visible from a quite large TEM01 mode at the dark port.

Here are the parameters I used:

MICH gain -10   PRCL gain -0.1

Normalization of both error signal on POP22_I with factor 0.004

Triggering on POP22: in at 100, out at 20 for both MICH and PRCL.

POP55 demodulation phase -9

MICH and PRCL control signal limits at 2000 counts

 

There is a high frequency (628 Hz) oscillation going on when locked (very annoying on the speakers...), but reducing the gain made the lock less stable. I could go down to MICH=-1.5 and PRCL=-0.02, still being able to acquire the lock. But the oscillation was still there. I suspect that it is not due to the loops, but maybe some resonance of the suspension or payload (violin mode?). There is still some room for fine tuning...

Lock is acquired without problems and maintained for minutes.

Have a nice week-end!

  8375   Fri Mar 29 19:23:49 2013 Gabriele, JenneFrogsLSCAnalog whitening filter of REFL55 not switching

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

The digital one is actually switching. We decided to keep the digital de-whitening on to compensate for the analog one. Otherwise we get a very bad shape of the PDH signal. Sorry Rana...

  8374   Fri Mar 29 17:24:43 2013 JamieUpdateComputersFB RAID power supply replaced

Steve ordered a replacement power supply for the FB JetStor power supply that failed a couple weeks ago.  I just installed it and it looks fine.

  8373   Fri Mar 29 16:35:54 2013 SteveUpdate40m UpgradingEndtable upgrade

 

 My full effort to get the optical table enclosure ready for the lab has failed today.

What I did: cut IR thermashield sheets to size for sides and top and sandwitched them beetween 1" and 1/4" acrylic.

                   The carpenter shop recut the bottom o -ring groove to 0.250" wide and 0.150" deep.

                   O-ring was cut and installed. So this was ready to go lab.

                  NO, I realized that the liner yellow acrylic was not cut correctly. It was larger than 1" wall.

                 The shop is cutting them down to fit and I have to resize IR shields

  8372   Fri Mar 29 16:30:27 2013 SteveUpdate40m Upgrading table leveling mounts go here

Quote:

Quote:

Steve's suggestion for how to level the end table using "swivel leveling mounts":

end-tabel-leveling.png

 

1, level table with 4 swivels, lock nuts  3/16 -vertical alignment    2, lock this position with 4 x 1/4-20 (these are in place already) to hold table in horizontal direction

   Adjustable mount is here to compensate for bad tilting  floor <http://www.mcmaster.com/#6111K52>  

      See grouting plans of the past   http://nodus.ligo.caltech.edu:8080/40m/7248

 

  8371   Fri Mar 29 00:27:06 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : Assembly

I am currently putting together all components so that they are ready to go on the table once leveling and installation of shield box is done. All dirty optics were drag wiped. These are stored in the cupboard along the Y arm.
I could not find the fiber coupling mount on the old endtable. Also the harmonic separator that reflects the trans beam to PDs and camera is labelled Y1-1064 (??) and I don't know what's the deal with this.
I am nearly 70% done with assembling…so the ex-endtable is almost empty.

Yet to do:
1. Mount 2" optics
2. Hunt, gather and mount appropriate lenses

Points I did not notice earlier:
We need some good 2" lens mounts and also order 2" lenses for IPANG and trans beam.

 

IMG_0433.JPGIMG_0434.JPG

  8370   Thu Mar 28 23:06:48 2013 AnnalisaUpdateAuxiliary locking"Alberto" NPRO laser again on PSL table

 "Alberto"NPRO laser has been moved again on PSL table in order to make a measurement of the beat note varying also the PSL temperature.

It is useful because if the PSL temperature would drift  we have to know which is the NPRO temperature that returns the beat.

I'm going to measure it tomorrow.

 

 

  8369   Thu Mar 28 23:00:30 2013 AnnalisaUpdateABSLBeat note of ATF auxiliary laser found

 

The beat note for the ATF lab laser has been found. 

The measurement has been carried out in the same way as described in elog 8368.

The only difference is that in this case I started from a temperature of 35.2 degC, and I reduced it until the minimum which was 30.71 degC. No beat note in this range.

Then I rised on the temperature and I found the first beat note at 41.46 degC. It has been detected at a frequency of about 120 MHz with an RF power of -53 dBm and a frequency fluctuation of about  +/- 5 MHz. 

I tried to improve the alignment to have a stronger beat, but it was the maximum I could reach. Maybe I could increase the power hitting the photodiode, which was 0.453 mW. 

 

 

  8368   Thu Mar 28 19:32:22 2013 AnnalisaSummaryAuxiliary lockingBeat note found!

 

 I plot the variation of the beat note frequency as a function of "Alberto" NPRO laser's temperature.

After some discussion, now I'm going to vary the PSL temperature and find the auxiliary NPRO temperature matching to have the beat note between the two.

  8367   Thu Mar 28 12:50:52 2013 JenneUpdateComputersc1lsc is fine

 Manasa told me that she did things in a different order than her old elog. 

She had

(1) ssh'ed to c1lsc and did a remote shutdown / restart,

(2) restarted fb,

(3) restarted the mxstream on c1lsc,

(4) restarted each model individually in some order that I forgot to ask.

However, with the situation as in her "before" screenshot, all that needed to be done was restart the mxstream process on c1lsc. 

Anyhow, when I looked at the OAF model, it was complaining of "no sync", so I restarted the model, and it came back up fine.  All is well again.

  8366   Thu Mar 28 10:44:30 2013 ManasaUpdateComputersc1lsc down

c1lsc was down this morning.

I restarted fb and c1lsc based on elog

Everything but c1oaf came back. I tried to restart c1oaf individually; but it didn't work.

Before:

cds_FE.png

After:

cds_fe1.png

  8365   Thu Mar 28 07:58:15 2013 SteveUpdateGreen LockingNPRO repair

Quote:

 

 JDSU can repair the Lightwave M126-1064-700 NPRO, sn 415  They do not need the Controller sn 516 

 Posted in the 40m Wiki_ PSL_ NPRO  cost repair and/or option to buy Innolight laser as replacement

 NPRO shipped out for evaluation yesterday under RMA 18022707

  8364   Thu Mar 28 04:21:40 2013 ranaUpdateLockingPRMI progress - even better!

 

 A key step was turning off the whitening filters. With the previous setting (G = 15 dB, white on), the error signals (post anti-whitening) had amplitudes of ~500 counts. This means that they can go as high as (150/15)^2 * 500 = 50000 counts on the ADC.

The purpose of the whitening filter is to match the noise / range of the signal to the ADC. What we would like to do is use the minimum gain so as to make the RFPD electronics noise + shot noise be ~equal to the ADC noise. i.e.

sqrt(V_PD^2 + v_shot^2) * G_white = V_ADC

The RFPD noise is ~3 nV before the internal preamp. The MAX4107 has a gain of 10. There is a factor of 1/2 from the voltage division of the RFPD's 50 Ohm series resistor and the input impedance of the mixer. There is also a power splitter between the PD output and the mixer which gives us a 3 dB loss. The mixer has a conversion loss of ~5-6 dB depending upon the LO level.

V_PD = 3e-9 * (10 * 1/2 * 1/sqrt(2) * 1/2) = 5e-9 V/rHz   (this is already bad; the signal coming out of the mixer needs to be amplified by x10 before going out to the whitening board).

In any case, its clear that we need something like 60 dB of gain for the PD noise to match the ADC noise. This is why increasing the whitening gain improves the error signal's SNR, reduces the hash driving the optics, and improves the locking. We should run with 45 dB gain and the switch on whitening after the lock.

 

Even better would be to modify the LT1128 input stage of the card to have the single stage of fixed whitening as we did for iLIGO. Then we can have triple whitening in lock.

  8363   Thu Mar 28 01:59:33 2013 JenneUpdateLockingPRMI progress - even better!

[Gabriele, Rana, Jenne, Jamie, Lisa, Zach]

We tweaked some things after dinner, and our locks got longer (~10sec) and more frequent!

What happened / notes:

* Increased analog gain from 15dB to 27dB for REFL55 I&Q.

* No analog whitening during lock acquisition.  (Need trigger + wait so whitening comes on after ~1sec...but this is not our limitation right now).

* Limit MICH and PRCL control to 5000, so that we don't kick optics too much, which makes them take too long to settle.

* ITMX and ITMY Vio2 filters turned off (PRM still has it on) in the SUS-optic_LSC module.

* MICH and PRCL DoFs triggering on POP22I, with levels 200 & 50.  FMs 4&5 always on.

* MICH and PRCL FM2 triggering on POP22I with levels 400 & 50.

* MICH gain = -0.200

* PRCL gain = +0.150

* MICH and PRCL normalization using POP22I, with matrix values 0.00160 .  This value is ~1/600, where 600 was the peak value of POP22I_ERR.

* REFL 55 phase set back to -15, to minimize PRCL signal in I phase.

* Checked signs for ITMX and ITMY in output matrix for MICH.  Lock MICH using only ITMX or ITMY, find sign to hold on the dark fringe for each.  +1 for ITMY, -1 for ITMX was correct.

* Tweaked up the oplev servos.  See separate elog 8362.  May need more tweaking, such as increasing the UGF, engaging 1Hz resonant gain.

* May need better coil actuator balancing on suspensions at 1Hz.

* Found a weird thing in DTT, which went away after closing and reopening, when looking at time series.  Sometimes we would see a square wave-like jump in the signals, all signals at the same time, with a frequency of 16.6Hz.  This was not present in other data retreival programs, like Jamie's getdata python script.

* We are not sure right now why we are falling out of lock.  We need to investigate more signals, to try to figure out what our current problem is.

* Reduced the amount of misalignment with the "misalign" script, to reduce hysteresis.

To Do / ideas:

* Calibrate oplev signals - see if one optic is moving more than others.

* Calibrate ERR and CTRL  - look at CTRL in meters, see if cavities are moving around like crazy.

* Calibrate POP22 using something like an AM laser modulation trick into units of PRCL SB gain. Compare with expectation - are we locked optimally, or do we have more power that we can be getting out?

* Try feeding back PRCL CTRL to MC2, to make the laser to follow the power recycling cavity, in hopes of reducing angular motion.  Rana tried this quickly with a 1 in the output matrix, but this kicked the MC out of lock - need to try smaller values.

 

  8362   Thu Mar 28 00:31:11 2013 JenneUpdateSUSPRMI optics' oplev servos tuned

[Rana, Gabriele, Jenne, Jamie, Lisa, Rana]

We have tuned the oplev servos for PRM, BS, ITMX, ITMY.  For each, we measured the servo transfer function.  Most had a UGF ~ 3Hz.  For those, we increased the gain by a factor of 2, and engaged the 3.3Hz resonant gains. The other case, such as PRM yaw, the gain was already okay, we just needed to engage the resonant gain.  We also checked the new phase margin, and for some of them switched the elliptic lowpass to 50Hz rather than 30 or 35.

Before and afters:

PRM_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

PRM_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

bs_pit.pdf

bs_yaw.pdf

ITMY_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

ITMY_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

itmx_pit1.pdf

itmx_yaw.pdf

We need to, as a last check, look at the spectra before and after to ensure that no modes (like bounce or roll) are newly excited.

  8361   Wed Mar 27 21:53:21 2013 AnnalisaUpdateABSLBeat note of ATF auxiliary laser

After measuring the beat note, the "Alberto" NPRO auxiliary laser has been moved from the PSL table to the POY table. Its beam profile is going to be measured. It's going to be used as green laser on the END table, in place of the broken one.

The auxiliary laser borrowed form ATF lab (which will be used for the ABSL measurement) has been set on the PSL table to make a measurement of the beat note between it and the main laser.

The setup is mostly the same of the previous beat note measurement . In this case, laser input power is 326 mW, so I needed to replace one of the mirrors of the steering optics with a BS 50% reflecting in order to have less than 1 mW on the PD.

Now, the total power on the PD is less than 0.5 mW.

I didn't measure the beat note yet to leave the PSL table as quite as possible for the locking procedures.

To do:

Measure the beat note, fiber coupling the NPRO laser to bring it to the POY table.

 

  8360   Wed Mar 27 15:46:13 2013 JenneUpdateLockingPRMI progress

[Gabriele, Jenne]

We have achieved PRMI locks of the order ~5 seconds! Here is an example lock:

PRMI_sb_27Mar2013.png

This was actuating MICH on the ITMs (+1 for ITMY, -1 for ITMX in the output matrix), and PRCL on PRM (+1). 

PRCL gain was +1, MICH gain was -10.

PRCL signal was normalized with POP22I with a matrix value of 0.003 . (No normalization of MICH).

Both PRCL and MICH were triggered on POP22I with high thresh of 200, low of 50.  MICH and PRCL FM2 (integrators) were triggered on POP22I with thresh of 400 and low thresh of 50.  FMs 4 and 5 were on for both MICH and PRCL always.

We zoomed in on the MICH_OUT signal, and the instability looks like it is around 300 Hz.  We aren't sure what this is.  I think this is a similar frequency to an oscillation that Yuta saw, but I'll have to check the old elogs.

PRM and BS SUS_LSC_POS filter banks both have notches between 1280-1290Hz.  The ITMs do not have this "Vio2" filter.

This is one of the first locks with the new triggering of both the INPUT and OUTPUT of the control filter banks.  I modified the lsc model before lunch. 

To do:  Where is this 300Hz coming from, and what can we do about it?  Why are we losing lock?  It's not due to the oscillation - maybe too much afternoon seismic?  Steve says he went next door and the rock monster / river is on medium/high.

 

  8359   Tue Mar 26 20:20:10 2013 CharlesUpdateISSISS Design Plans - Servo Noise Analysis

In order to allow other individuals besides myself to consider the proposed design of the ISS, I have created a publicly available CircuitLab drawing, which can be found here: CircuitLab Drawing. For simplicity, I have used ideal op-amps without voltage rails or their associated power supplies. In the actual implementation of the ISS, we will most likely also have trim resistors to ensure a zero offset for each op-amp. We interpret the PD as a voltage source for simplicity and I will use an actual summing amplifier in place of the summing junction used in the diagram.

The diagram linked above is simply a naive copy of a design by Rich Abbott so there are most likely mistakes and/or unnecessary elements, but it is a work in progress. I began discussing, with Jamie, the relative use of the first few filter stages in the servo. As far as my understanding goes, the first 'stage' was part of cascade of op-amps that served to convert a differential input from the PD into a single DC signal referenced to ground. Indeed, the first stage of my diagram (U1) is simply a unity-gain low-pass filter with f~5 MHz. Additionally, the second filter 'stage', U2, is also a unity-gain low-pass filter although it introduces a phase shift of 180 deg as the input to the second stage is on the inverting input of the op-amp. These characteristics were determined using LISO and examining the transfer function.

Noise analysis was also performed for the above circuit. The noise from various elements is examined at the output of the servo (labeled as 'outU6' in my LISO file). In the attached diagram, we see the voltage noise at the output from each op-amp as well as the sum of all the various noises, which includes resistor noise and current noise from the inputs of each op-amp. These are LISO's standard considerations and it is also worthwhile to note that the result is not referred to the circuit input, but as we have the transfer function of the whole servo, referring the noise to the input is trivial.

I have also included the following output for the sake of completeness.

from 1 Hz onwards noise by OP:I+ (U3) dominates.

from 38.6812 Hz onwards noise by R(R24) dominates.

from 115.478 Hz onwards noise by R(R11) dominates.

 

 

  8358   Tue Mar 26 17:32:30 2013 ChloeUpdate QPD/ECDL Progress

I built the summing/subtracting circuit on the breadboard, and hooked this up with one of the other QPDs I found (image of setup attached). I wasn't able to get this to read the correct signals when testing with a laser pointer after a couple of hours of troubleshooting... I will hopefully get this working in the next day or 2...

I'm going to read up on ECDL stuff for Tara tonight and hopefully figure out what sort of laser diode we should purchase, since I'm meeting with Tara tomorrow. experimenting

  8357   Tue Mar 26 17:27:17 2013 KojiUpdateIOOTuning MC length/FSR

Please add the discussions on the cavity absolute length (and its change, adjustment precision),
identification of the peaks, before/after comparison of the plot, the effect of the MC REFL PD response,
and comparison of the cavity linewidth vs deviation of the 55MHz SBs from the resonance.

  8356   Tue Mar 26 16:36:24 2013 SteveUpdate40m Upgrading table leveling mounts go here

Quote:

Steve's suggestion for how to level the end table using "swivel leveling mounts":

end-tabel-leveling.png

 

1, level table with 4 swivels, lock nuts  3/16 -vertical alignment    2, lock this position with 4 x 1/4-20 (these are in place already) to hold table in horizontal direction

   Adjustable mount is here to compensate for bad tilting  floor <http://www.mcmaster.com/#6111K52>  

      See grouting plans of the past   http://nodus.ligo.caltech.edu:8080/40m/7248

ELOG V3.1.3-